Paige is considering job offers from two companies. Company A offered her a starting salary of $55,000 with a $2100 raise at the end of each year. Company B offered her a starting salary of $55,000 with a 3.6% raise at the end of each year.

Let f(t)f(t) represent Paige's salary at Company A tt years after accepting a position at Company A, and let g(t)g(t) represent Paige's salary at Company B tt years after accepting a position at Company B. Complete the table of values below.

tt f(t)f(t) g(t)g(t)
0 55000 55000
1 56980
2 58960
3
Answer #1

Auser Loubing A- (t)55,00o+ alo0t (t55,000+ 2100 K f(X)= 55,000 flo) = 55,000+2,00 When t=2 55,000+2100x2 f(+)= f(t): 55,000+whenた 2 (4) = 55,Doo(Ms) g(4) 59,031.23 2(2) 二 when t 3 (4)= 61,156, Y061 ,(3) (大) 0 55, 000 55000 57,100 56,980 2 59,200 59,